OpenStudy (anonymous):

Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)? A) g(x) − f(x) = 3x − 8 − 5x B) g(x) − f(x) = 3x − 10 − 5x C) g(x) − f(x) = −2x − 8 − 3x D) g(x) − f(x) = −2x − 10 − 3x I got C.. am i right?

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

expanded, g(x) - f(x) would be 3x - 9 -5x +1, condensed would be -2x -8

OpenStudy (anonymous):

so i was wrong?

OpenStudy (anonymous):

yes, A is the closest by expanding the -2x to the 3x-5x

OpenStudy (anonymous):

that was the other one i thought it might be, thanks!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

it wasnt a...

OpenStudy (anonymous):

what are you going off of to determine the answer?

OpenStudy (aum):

They simplified the constant terms but left the x terms unsimplified: g(x) − f(x) = 3x − 9 - (5x − 1) = 3x - 9 - 5x + 1 = 3x - 9 + 1 -5x = 3x - 8 - 5x So A.